OpenStudy (thats_kyy):

A set of equations is given below: Equation A: y = x + 1 Equation B: y = 4x + 5 Which of the following steps can be used to find the solution to the set of equations?

OpenStudy (thats_kyy):

- x + 1 = 4x + 5 - x= 4x + 5 - x + 1 = 4x - x + 5 = 4x + 1

OpenStudy (thats_kyy):

@perl

OpenStudy (thats_kyy):

@Ashleyisakitty

OpenStudy (thats_kyy):

Someone Help Me!! Please

OpenStudy (thats_kyy):

@campbell_st Help Me Please

OpenStudy (campbell_st):

both equations equal y, so then you can equate the 2 equations...

OpenStudy (thats_kyy):

@campbell_st So Is The Answer B

OpenStudy (thats_kyy):

@skittles_for_life6422 Do You Think You Can Help Me? With This Question

OpenStudy (skittles_for_life6422):

sure

OpenStudy (thats_kyy):

okay so do you see the question and answer choices i put down @skittles_for_life6422

OpenStudy (skittles_for_life6422):

yeah i think it is a

OpenStudy (thats_kyy):

okay thank you @skittles_for_life6422